# Continuum-Field-Theoretical Construction of Conserved Currents, Non-Invertible Symmetries, and Mixed Anomalies in (3+1)D Non-Abelian Topological Order

### v2

@@ Abstract
A modern perspective views topological order as a symmetry-breaking phase of generalized symmetries, making the identification of such symmetries, especially in non-Abelian topological orders, a fundamental task. In this work, we investigate generalized symmetries, with particular emphasis on non-invertible ones, in three-dimensional non-Abelian topological orders hosting both particle- and loop-like excitations. We adopt a continuum topological field theory description, focusing on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twisted term. This field theory supports Borromean-Rings braiding and realizes non-Abelian topological order, which for $G=(\mathbb{Z}_{2})^{3}$ admits a microscopic realization via the $\mathbb{D}_{4}$ Kitaev quantum double lattice model. We systematically identify all generalized symmetry operators by extracting conserved currents from the equations of motion. Two distinct classes of currents emerge: type-I currents, which generate invertible higher-form symmetries, and type-II currents, which give rise to non-invertible higher-form symmetries. The non-invertibility originates from projectors accompanying the symmetry operators, which restrict admissible gauge-field configurations. We further analyze the fusion rules of these symmetries, showing that invertible symmetries admit inverses, while non-invertible symmetries fuse through multiple channels. Finally, we study mixed anomalies among these generalized symmetries by simultaneously coupling multiple currents to proper types of background gauge fields and examining their gaugeability. We identify two types of mixed anomalies: one cancellable by topological field theories in one higher dimension, and another representing an intrinsic gauging obstruction encoded in the (3+1)(3+1)D continuum topological field theory. Our results provide a concrete field-theoretical construction of conserved currents and a unified characterization of invertible and non-invertible higher-form symmetries in three-dimensional non-Abelian topological orders, offering a natural framework for studying (2+1)(2+1)D systems via symmetry topological field theory and topological holography.
@@ I Introduction
Symmetry has long been a central organizing principle in physics. Broadly speaking, a symmetry of a physical system is a transformation that leaves the underlying theory invariant when applied to the system as a whole. A familiar example is provided by the one-dimensional [^1] In this paper, “(n+1)(n+1)D” refers to $(n+1)$-dimensional spacetime with $n$-dimensional real space. We avoid using “nnD” or “(n)(n)D” unless otherwise specified. When referring specifically to spatial dimensions or to the dimensions of geometric objects, we use “nnd”, e.g., 33d ground state, 33d topological order, 22d square lattice, and 11d submanifold. (11d) transverse-field Ising model, which is invariant under a simultaneous flip of all spins. This global spin-flip operation acts on the entire system and leaves the Hamiltonian unchanged, and is therefore identified as a symmetry. Local operators may transform nontrivially under this operation and are said to carry symmetry charge. Since performing the global spin flip twice returns the system to itself, this symmetry has a $\mathbb{Z}_{2}$ group structure.
@@
In recent years, the notion of symmetry has been substantially generalized from multiple perspectives across condensed matter, high-energy, and mathematical physics [1, 2, 3, 4, 5, 6, 7, 8]. Two complementary aspects are particularly important. First, symmetry operations need not act on the entire system; instead, they may be supported on submanifolds [1]. As a result, charged operators become extended rather than point-like, leading to the concept of higher-form symmetry. Within this framework, the $\mathbb{Z}_{2}$ symmetry of the 11d transverse-field Ising model discussed above is an example of a 0-form symmetry.
@@
A canonical illustration of higher-form symmetry is provided by the 22d toric code model [9] on a square lattice, where symmetry operators can be constructed on closed loops rather than on the full lattice. These loop operators commute with the Hamiltonian and generate symmetries associated with one-dimensional submanifolds embedded in the two-dimensional system. Operators supported on one-dimensional paths that intersect these loops transform nontrivially under the symmetry and are therefore charged. Such symmetries are referred to as $1$-form symmetries, reflecting the co-dimensionality of their support in 22d real space. At long wavelengths, the supports of higher-form symmetries are manifold-like and can be continuously deformed without affecting their action. There exists another important generalization, namely subsystem symmetries, whose supports are rigid rather than deformable and are typically fixed to specific geometric patterns such as lines, fractals, or more exotic structures [10, 11, 12, 13]. In this work, however, we focus exclusively on higher-form symmetries and leave subsystem symmetries for future investigation.
@@
Second, the algebraic structure organizing symmetry operators need not be a group [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. More generally, it may be described using the language of category theory, giving rise to non-invertible or categorical symmetries. A representative example arises in the 11d transverse-field Ising model at criticality, which admits a symmetry operation implementing Kramers-Wannier duality. This operation exchanges spin and domain-wall variables and commutes with the Hamiltonian, yet it does not admit an inverse in the conventional sense: acting with it projects out a subset of states. Such an operation therefore realizes a non-invertible symmetry. For non-invertible symmetries, composing two symmetry operations may not yield a single operator but rather a linear combination of symmetry operators governed by fusion rules. By contrast, for invertible symmetries the fusion rules reduce to the familiar group composition law.
@@
These modern generalizations of symmetry provide powerful tools for the study of quantum many-body systems, particularly topologically ordered phases. Such phases, traditionally viewed as lying beyond the Landau-Ginzburg paradigm, can be reinterpreted as phases that spontaneously break generalized symmetries. For example, the 22d toric code model on a torus exhibits a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 11-form symmetry generated by operators supported on the two noncontractible cycles of the torus. Since these operators do not commute, no ground state can remain invariant under both simultaneously, leading inevitably to topological ground-state degeneracy. This phenomenon can be understood as spontaneous breaking of a higher-form symmetry. In this way, generalized symmetries offer a unifying framework for understanding both the structure and robustness of topologically ordered phases.
@@
Building on the concept of generalized symmetry, a framework known as symmetry topological field theory (SymTFT), also referred to as the symmetry-topological-order correspondence, has been rapidly developed. The central idea is that a $(d+1)$-dimensional topological quantum field theory (TQFT), or equivalently a (d+1)(d+1)D topological order, encodes the symmetries of a (d1+1)(d-1+1)-dimensional quantum field theory. In this correspondence, topological operators in the bulk encode generalized symmetries of the TQFT and can terminate on the boundary in different ways, thereby characterizing distinct lower-dimensional systems. This framework has found broad applications, including the classification of gapped and gapless phases, the diagnosis of generalized symmetry anomalies, the analysis of renormalization-group flows, the study of mixed-state phases and phase transitions, and so on [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. To date, SymTFT has been particularly successful for (2+1)(2+1)D bulks and (1+1)(1+1)D boundaries. Extending this framework to (2+1)(2+1)D systems therefore naturally motivates a detailed understanding of generalized symmetries in (3+1)(3+1)D topologically ordered phases.
@@
As effective field theories of topological order, TQFTs are expected to encode generalized symmetries in a natural and systematic manner. For instance, Wilson operators in Chern-Simons theory [36, 37, 38] act as generators of $1$-form symmetries. Analogous to the role of Chern-Simons theory in describing 22d topological orders, $BF$ theory [39, 40], together with its twisted variants [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], provides a unified field-theoretical framework for characterizing 33d topological orders and gauged symmetry-protected topological phases with particle and loop excitations. Geometrically, the canonical $BF$ term bdab\,da (with $b$ a $2$-form field and $a$ a $1$-form field) captures the characteristic particle-loop braiding in 33d topological order. [^2] Throughout this paper, we adopt a condensed notation for differential forms and omit the wedge symbol $\wedge$ when no confusion arises. For example, bdab\,da stands for $b\wedge da$, and similarly aadaaada denotes topological terms of $a\wedge a\wedge da$ type.
@@
More intricate braiding processes, including multi-loop and Borromean rings braiding, are encoded by twisted terms such as aadaaada, aaaaaaaa, and aabaab [40, 44, 48, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. These braiding processes are subject to compatibility conditions imposed by gauge invariance [49]. Additional terms, such as the twisted $bb$ term and the $\theta$ term dadada\,da, encode emergent fermionic statistics [45, 50, 66, 67] and topological responses [41, 43, 51, 68, 69, 70, 71, 72], respectively. Further developments revealed symmetry fractionalization on loop excitations, leading to mixed three-loop braiding and the classification of symmetry-enriched topological phases in higher dimensions [73, 74, 75]. The fusion and shrinking rules of topological excitations in 33d topological orders have also been investigated within the $BF$-theory framework [76]. Motivated by this field-theoretical foundation, a diagrammatic representation of higher-dimensional topological orders, including 44d cases [77], was recently developed in Ref. [78]. Given this progress, a natural question arises: how can generalized symmetries of 33d topological orders be systematically identified directly from effective field theory?
@@
In this work, we address this question by studying generalized symmetries in a class of (3+1)(3+1)D topological quantum field theories describing three-dimensional non-Abelian topological orders with nontrivial Borromean-Rings braiding. We focus on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twisted term [48, 49, 76]. For the special case $G=(\mathbb{Z}_{2})^{3}$, the resulting non-Abelian topological order admits a microscopic realization in terms of the $\mathbb{D}_{4}$ Kitaev quantum double lattice model [79]. These field theories therefore provide a concrete continuum description in which higher-form generalized symmetries, including non-invertible ones, can be systematically analyzed. In quantum mechanics, an operator commuting with the Hamiltonian signals a symmetry, while in field theory the corresponding notion is a conserved current. Guided by this principle, we identify generalized symmetries directly from the effective field theory by extracting conserved currents from the equations of motion of the gauge fields.
@@
We find that the equations of motion naturally separate into two distinct classes, leading to two types of conserved currents. Type-I currents are conserved identically and generate invertible higher-form symmetries. By contrast, type-II currents are conserved only upon imposing additional constraints on the gauge-field configurations. As a result, the associated symmetry operators necessarily involve projector structures that enforce these constraints. Since such projectors do not admit inverses, the corresponding higher-form symmetries are intrinsically non-invertible.
With the explicit symmetry operators constructed from these conserved currents, we analyze their fusion rules by examining the composition of symmetry operations. We show that invertible higher-form symmetries possess well-defined inverses and obey group-like fusion rules, whereas non-invertible symmetries fuse through multiple channels, reflecting the presence of projector-induced constraints. Finally, we investigate mixed anomalies among these generalized symmetries by simultaneously coupling multiple conserved currents to appropriate background gauge fields and attempting to gauge the corresponding symmetries. We identify two distinct types of mixed anomalies: one that can be canceled by embedding the theory into a topological field theory in one higher dimension, and another that represents an intrinsic gauging obstruction already encoded in the (3+1)(3+1)D continuum topological field theory.
@@
This paper is organized as follows. In Sec. II, we review generalized symmetries, emphasizing the roles of conserved currents and topological operators. In Sec. III, we study generalized symmetries in the effective field theory of a class of 33d non-Abelian topological orders and construct the corresponding symmetry operators from conserved currents. Sec.p IV is devoted to the analysis of fusion rules. In Sec. V, we discuss anomalies of generalized symmetries. We conclude in Sec. VI with a summary and an outlook.

### v3

# Non-invertible symmetries and mixed anomalies from conserved current construction in (3+1)D twisted $BF$ topological quantum field theories

@@ Abstract
We develop a current-based construction of generalized symmetries in (3+1)(3+1)D twisted $BF$ topological quantum field theories (TQFTs), focusing on intrinsically non-invertible higher-form symmetries and their mixed anomalies. Starting from the equations of motion, we extract conserved currents and exponentiate the corresponding charges to obtain topological symmetry operators. This gives a step-by-step procedure for constructing symmetry operators, fusion, and anomaly diagnostics directly from the continuum action. We focus on twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twist, where $a$’s and $b$ are 1-form and 2-form gauge fields, respectively. These theories realize non-Abelian (3+1)(3+1)D TQFTs supporting Borromean-rings braiding and describe three-dimensional non-Abelian topological orders in condensed matter. For $G=(\mathbb{Z}_{2})^{3}$, a microscopic realization is given by the $\mathbb{D}_{4}$ Kitaev quantum double model. Two distinct classes of conserved currents emerge: Type-I currents generate invertible higher-form symmetries with group-like fusion, while Type-II currents require additional consistency conditions on gauge-field configurations, leading to intrinsically non-invertible symmetries dressed by projectors. We compute the fusion algebra: invertible operators admit inverses, while non-invertible ones exhibit multi-channel fusion governed by projector fusion. We diagnose mixed anomalies by coupling multiple conserved currents to background gauge fields, revealing two outcomes: anomalies canceled by anomaly inflow from a higher-dimensional theory, and intrinsic gauging obstructions encoded in the (3+1)(3+1)D continuum theory. Overall, our results provide a unified and practical approach for constructing and characterizing higher-form symmetries, which can be extended to more general TQFTs and topological orders.
@@ I Introduction
Symmetry has long been a central organizing principle in physics. Broadly speaking, a symmetry of a physical system is a transformation that leaves the underlying theory invariant when applied to the system as a whole; equivalently, in the operator language it is implemented by an operator commuting with the Hamiltonian (or with the dynamics in the continuum). A familiar example is provided by the one-dimensional [^1] In this paper, “(n+1)(n+1)D” refers to $(n+1)$-dimensional spacetime with $n$-dimensional real space. We avoid using “nnD” or “(n)(n)D” unless otherwise specified. When referring specifically to spatial dimensions or to the dimensions of geometric objects, we use “nnd”, e.g., 33d ground state, 33d topological order, and 22d square lattice. (11d) transverse-field Ising model, which is invariant under a simultaneous flip of all spins. This global spin-flip operation acts on the entire system and leaves the Hamiltonian unchanged, and is therefore identified as a symmetry. Local operators may transform nontrivially under this operation and are said to carry symmetry charge. Since performing the global spin flip twice returns the system to itself, this symmetry is invertible and has a $\mathbb{Z}_{2}$ group structure.
@@
In recent years, the notion of symmetry has been substantially generalized from multiple perspectives across condensed matter, high-energy, and mathematical physics [1, 2, 3, 4, 5, 6, 7, 8]. Two complementary aspects are particularly important. First, symmetry operations need not act on the entire system; instead, they may be supported on submanifolds [1]. As a result, symmetry operators can be extended (line/surface) operators, and charged operators become extended rather than point-like, leading to the concept of higher-form symmetry. Within this framework, the $\mathbb{Z}_{2}$ symmetry of the 11d transverse-field Ising model discussed above is an example of a 0-form symmetry. Second, the algebraic structure organizing symmetry operators need not be a group [9, 10, 11, 12, 13, 14, 15, 7, 16, 17]. More generally, symmetry is encoded by the fusion of topological operators/defects, and is naturally formulated using the language of category theory, giving rise to non-invertible or categorical symmetries.
@@
As an illustration of the first aspect, the 22d toric code model [18] on a square lattice admits symmetry operators supported on closed loops rather than on the full lattice. In condensed-matter language, these loop operators commute with the Hamiltonian and act within the topological ground-state sector; in the continuum/TQFT language, they are topological line operators (Wilson loops) generating higher-form symmetries. Operators supported on one-dimensional paths that intersect these loops transform nontrivially under the symmetry and are therefore charged. Such symmetries are referred to as $1$-form symmetries, reflecting the co-dimensionality of their support in 22d real space. At long wavelengths, higher-form symmetry operators are topological in the sense that their supports are manifold-like and can be continuously deformed without affecting their action. There exists another important generalization, namely subsystem symmetries, whose supports are rigid rather than deformable and are typically fixed to specific geometric patterns such as lines, fractals, or more exotic structures [19, 20, 21, 22]. In this work, however, we focus exclusively on higher-form symmetries and leave subsystem symmetries for future investigation.
@@
As an illustration of the second aspect, a representative example of a non-invertible symmetry arises in the 11d transverse-field Ising model at criticality, which admits a symmetry operation implementing Kramers–Wannier duality. This operation exchanges spin and domain-wall variables and commutes with the Hamiltonian, yet it does not admit an inverse in the conventional sense: acting with it projects out a subset of states. Such an operation therefore realizes a non-invertible symmetry. For non-invertible symmetries, composing two symmetry operations may not yield a single operator but rather a linear combination of symmetry operators governed by fusion rules. By contrast, for invertible symmetries the fusion rules reduce to the familiar group composition law.
@@
These modern generalizations of symmetry provide powerful tools for the study of quantum many-body systems, particularly topologically ordered phases. Such phases, traditionally viewed as lying beyond the Landau–Ginzburg paradigm, can be reinterpreted as phases characterized by generalized-symmetry structures, and are often described as exhibiting spontaneous breaking of generalized symmetries in the topological ground-state sector. For example, the 22d toric code model on a torus exhibits a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 11-form symmetry generated by loop operators supported on the two noncontractible cycles of the torus. Since these operators do not commute, no ground state can remain invariant under both simultaneously, leading inevitably to topological ground-state degeneracy. This phenomenon can be viewed as a manifestation of higher-form symmetry breaking in the ground-state subspace. In this way, generalized symmetries offer a unifying framework for understanding both the structure and robustness of topologically ordered phases.
@@
Building on the concept of generalized symmetry, a framework known as symmetry topological field theory (SymTFT) has been rapidly developed. The central idea is that a $(d+1)$-dimensional topological quantum field theory (TQFT), or equivalently a (d+1)(d+1)D topological order in condensed-matter language, encodes the generalized symmetries of a (d1+1)(d-1+1)-dimensional quantum field theory. In this correspondence, topological operators/defects in the bulk encode generalized symmetries and can terminate on the boundary in different ways, thereby characterizing distinct lower-dimensional systems. This framework has found broad applications, including the classification of gapped and gapless phases, the diagnosis of generalized-symmetry anomalies, the analysis of renormalization-group flows, the study of mixed-state phases and phase transitions, and so on [15, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. To date, SymTFT has been particularly successful for (2+1)(2+1)D bulks and (1+1)(1+1)D boundaries. Extending this framework to (2+1)(2+1)D boundary systems therefore naturally motivates a detailed understanding of generalized symmetries realized in (3+1)(3+1)D TQFTs/topologically ordered phases.
@@
As effective field theories of topological order, TQFTs are expected to encode generalized symmetries in a natural and systematic manner. For instance, Wilson operators in Chern–Simons theory [34, 35, 36] act as generators of $1$-form symmetries. Analogous to the role of Chern–Simons theory in describing 22d topological orders, $BF$ theory [37, 38], together with its twisted variants [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], provides a unified field-theoretical framework for characterizing 33d topological orders and gauged symmetry-protected topological phases with particle and loop excitations. Geometrically, the canonical $BF$ term bdab\,da (with $b$ a $2$-form field and $a$ a $1$-form field) captures the characteristic particle–loop braiding in 33d topological order. [^2] Throughout this paper, we adopt a condensed notation for differential forms and omit the wedge symbol $\wedge$ when no confusion arises. For example, bdab\,da stands for $b\wedge da$, and similarly aadaaada denotes topological terms of $a\wedge a\wedge da$ type.
@@
More intricate braiding processes, including multi-loop and Borromean-Rings braiding, are encoded by twisted terms such as aadaaada, aaaaaaaa, and aabaab [38, 42, 46, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. These braiding processes are subject to compatibility conditions imposed by gauge invariance [47]. Additional terms, such as the twisted $bb$ term and the $\theta$ term dadada\,da, encode emergent fermionic statistics [43, 48, 64, 65] and topological responses [39, 41, 49, 66, 67, 68, 69, 70], respectively. Further developments revealed symmetry fractionalization on loop excitations, leading to mixed three-loop braiding and the classification of symmetry-enriched topological phases in higher dimensions [71, 72, 73]. The fusion and shrinking rules of topological excitations in 33d topological orders have also been investigated within the $BF$-theory framework [74]. Motivated by this field-theoretical foundation, a diagrammatic representation of higher-dimensional topological orders, including 44d cases [75], was recently developed in Ref. [76]. Given this progress, a natural question arises: how can generalized symmetries of 33d topological orders—equivalently, of their (3+1)(3+1)D effective TQFT descriptions—be systematically identified directly from continuum field theory, together with their fusion and anomaly data?
Reader’s guide (condensed-matter vs. TQFT language).
To minimize translation cost across communities, we will freely use both terminologies. In particular, “33d topological order” refers to the gapped topological sector captured by a (3+1)(3+1)D TQFT; point/loop excitations in the many-body system correspond to line/surface operators (topological defects) in the TQFT; and generalized symmetries are implemented by topological operators supported on submanifolds, whose fusion encodes the symmetry algebra. When we speak of “gauging” a generalized symmetry, we mean coupling to background gauge fields and testing gaugeability (possibly after adding local counterterms), thereby diagnosing potential ’t Hooft anomalies.
This work: a constructive and interdisciplinary route.
In this work we provide a fully constructive, model-resolved derivation of generalized symmetries in a nontrivial family of (3+1)(3+1)D TQFTs/topological orders. Our central output is a current-to-operator pipeline: starting from the action and equations of motion, we extract conserved currents and explicitly construct the associated topological symmetry operators/defects, together with their fusion rules and (anomalous) gaugeability. The same construction is intended to be useful from both sides of the condensed-matter/HEP interface: it offers a practical way to compute generalized-symmetry data of concrete three-dimensional topological orders described by twisted $BF$-type effective theories, and it organizes that data directly in the defect/anomaly language natural to continuum TQFT.
@@
Concretely, we study twisted $BF$ theories with gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$ and an $a\wedge a\wedge b$ twist [46, 47, 74], which realize non-Abelian (3+1)(3+1)D TQFTs supporting Borromean-Rings braiding and serve as effective continuum descriptions of certain three-dimensional non-Abelian topological orders. For the special case $G=(\mathbb{Z}_{2})^{3}$, the corresponding non-Abelian topological order admits a microscopic realization in the $\mathbb{D}_{4}$ Kitaev quantum double lattice model [77]. This setting provides a concrete arena in which higher-form generalized symmetries—including intrinsically non-invertible ones—can be identified and analyzed systematically.
Our approach is current-based. In quantum mechanics, an operator commuting with the Hamiltonian signals a symmetry; in continuum field theory, the corresponding diagnostic is a conserved current. Guided by this principle, we derive the relevant continuity equations from the equations of motion and use the resulting conserved currents to construct symmetry operators/defects. In this way, the generalized-symmetry data are obtained directly from conservation laws, and the same consistency conditions that make the currents well-defined determine the resulting fusion structure and (anomalous) gaugeability.
Main results and roadmap.
Our results can be summarized as a concrete route from continuum equations to symmetry data of (3+1)(3+1)D TQFTs/topological orders:
@@
  • Two classes of conserved currents from the equations of motion. We show that the equations of motion in the $a\wedge a\wedge b$ twisted $BF$ theories naturally separate into two distinct classes, leading to two qualitatively different types of conserved currents. Type-I currents are conserved identically (in the sense of Bianchi/Noether-type identities) and generate invertible higher-form symmetries. Type-II currents become conserved only after imposing additional constraints on admissible gauge-field configurations; these constraints are intrinsic to the twisted $BF$ structure and are required for the corresponding conserved quantities to be well-defined.
  • Explicit symmetry operators, including intrinsically non-invertible higher-form symmetries. From the type-I currents we construct standard topological symmetry operators with group-like composition, which in physical terms are the continuum counterparts of Wilson operators for particle/loop excitations. From the type-II currents we construct symmetry operators that must be dressed by projectors enforcing the above constraints. Because these projectors do not admit inverses, the resulting higher-form symmetries are intrinsically non-invertible. In particular, non-invertibility in our setting is not postulated: it is enforced by the consistency conditions behind the conservation laws.
  • Fusion algebra: group-like fusion vs. multi-channel fusion controlled by projector fusion. With explicit operators at hand, we compute the fusion rules by composing symmetry operators supported on the same submanifold (equivalently, by bringing the corresponding defects together). Invertible higher-form symmetry operators admit inverses and obey group-like fusion rules. By contrast, non-invertible higher-form symmetry operators exhibit genuinely non-group-like behavior: the product decomposes into multiple channels, and the channel structure is dictated by the fusion of the projector factors. We present representative examples that illustrate how the projector constraints reorganize the fusion algebra in a systematic, computable way.
  • Mixed anomalies and gauging: inflow-cancelable anomalies vs. intrinsic gauging obstructions. We diagnose anomalies by coupling multiple conserved currents to appropriate background gauge fields and testing gaugeability. This reveals mixed anomalies among the generalized symmetries (including between invertible and non-invertible ones). We further identify two qualitatively different outcomes: (i) anomalies that can be canceled by embedding into a topological field theory in one higher dimension (anomaly inflow), and (ii) intrinsic gauging obstructions already encoded in the (3+1)(3+1)D continuum theory. We illustrate these diagnostics both in familiar settings (e.g., the 33d $\mathbb{Z}_{2}$ topological order) and in the Borromean-Rings topological order described by the twisted $BF$ theories.
This structure makes the paper self-contained from a field-theory viewpoint: starting from the action and equations of motion, we extract conserved currents, construct the corresponding topological symmetry operators/defects, compute their fusion algebra, and finally analyze their (anomalous) gaugeability.
@@
This paper is organized as follows. In Sec. II, we review generalized symmetries, emphasizing the roles of conserved currents and topological operators. In Sec. III, we study generalized symmetries in the effective field theory description of a class of 33d non-Abelian topological orders and construct the corresponding symmetry operators from conserved currents. Sec. IV is devoted to the analysis of fusion rules. In Sec. V, we discuss anomalies of generalized symmetries. We conclude in Sec. VI with a summary and an outlook.